# DLF: Extreme Image Compression with Dual-generative Latent Fusion

### v1

@@ Abstract
Recent studies in extreme image compression have achieved remarkable performance by compressing the tokens from generative tokenizers. However, these methods often prioritize clustering common semantics within the dataset, while overlooking the diverse details of individual objects. Consequently, this results in suboptimal reconstruction fidelity, especially at low bitrates. To address this issue, we introduce a Dual-generative Latent Fusion (DLF) paradigm. DLF decomposes the latent into semantic and detail elements, compressing them through two distinct branches. The semantic branch clusters high-level information into compact tokens, while the detail branch encodes perceptually critical details to enhance the overall fidelity. Additionally, we propose a cross-branch interactive design to reduce redundancy between the two branches, thereby minimizing the overall bit cost. Experimental results demonstrate the impressive reconstruction quality of DLF even below 0.01 bits per pixel (bpp). On the CLIC2020 test set, our method achieves bitrate savings of up to 27.93% on LPIPS and 53.55% on DISTS compared to MS-ILLM. Furthermore, DLF surpasses recent diffusion-based codecs in visual fidelity while maintaining a comparable level of generative realism. Code will be available later.
@@ 1 Introduction
As image data continues to grow, achieving efficient compression while maintaining human perceptual quality has become crucial. At low bitrates, both traditional codecs [6, 10] and learning-based codecs [9, 17, 27] (optimized for PSNR or MS-SSIM [42]) often produce blurry reconstructions due to substantial information degradation [5]. In this context, generative image codecs [2, 32, 35], which enhance subjective quality by adding high-frequency details to reconstructed images [5], have attracted significant attention. To achieve effective generative coding at extremely low bitrates, recent studies [31, 43, 16] have employed generative visual tokenizers [41, 12] as their foundational approaches. These tokenizers cluster common semantic contents across the entire dataset into a compact codebook while simultaneously learning a powerful pixel generator based on the clustered information. As illustrated in Fig. 1 (a), if you do not look closely, many common semantic contents within the dataset can be reconstructed with acceptable quality using a small codebook. Consequently, such codecs [2, 32, 35] only need to compress and transmit a few token indices, thereby significantly enhancing compression ratio.
@@
However, when we zoom in each image to focus on the details of individual objects, we observe that the limited codebook capacity fails to capture these details accurately when compared with the ground truth. This inadequacy is particularly unfavorable for objects with precise geometric features, leading to a noticeable degradation in fidelity even without a reference for comparison. As illustrated in Fig. 1 (a), a reduced number of clusters results in severe distortion on a dog’s face. This observation underscores the limitation of the aforementioned codecs [2, 32, 35]: their backend tokenizer employs a single codebook to learn uniform latent clusters for all image contents, failing to balance the demand between the common semantic learning across the dataset and the diverse detail representation of the individual objects. Therefore, the question arises: How can we compress the latent in a more flexible manner, dynamically adapting to the content variations of each image?
@@
To tackle this challenge, we introduce a novel extreme image coding paradigm named Dual-generative Latent Fusion (DLF). As illustrated in Fig. 3, DLF decomposes the image latent into semantic and detail components, employing two parallel branches to compress each component according to their demands. The semantic branch inherits the clustering capability of the generative tokenizer, capturing the common content at the dataset-level with few bits. Meanwhile, the detail branch complements the shortcomings of the single codebook tokenizer, representing the diverse details through a large quantization space. By training with a rate-distortion loss, the detail branch learns to adaptively allocate more bits to the most critical regions for enhanced fidelity, meanwhile avoiding the excessive data overhead. Finally, the decoded semantic and detail features are fused using a latent adaptor, aligning them into the same latent space for high quality reconstruction using the pixel generator. As shown in Fig. 1 (b), our DLF not only reduces bit cost, but also significantly enhances the detail fidelity while preserving the overall perceptual quality.
@@
As adopting a dual-branch coding design, DLF will generate two bitstreams, similar to HybridFlow [30]. To minimize the overall bit cost, it is crucial to reduce redundancy between the two bitstreams. To this end, we introduce a cross-branch interaction mechanism that employs multiple attention-based layers to jointly process the intermediate features of both branches. This interaction not only optimizes bit allocation between the bottlenecks but also enables each branch to only retain the complementary information, thereby reducing redundancy. By contrast, HybridFlow adopts two independent encoders without joint design, significantly increasing bit cost due to redundancy.
@@
Powered by our advanced designs, our DLF achieves state-of-the-art (SOTA) compression performance across various datasets and metrics, especially at extremely low bitrates, such as 0.01 bpp. Qualitative results in Fig. 2 highlight the superiority of our method in terms of both generative realism and reconstruction fidelity. Compared to MS-ILLM [35], our approach achieves bitrate savings of 27.93% for LPIPS and 53.55% for DISTS on the CLIC2020 dataset. Additionally, our DLF significantly surpasses the recent diffusion-based codecs [7, 24, 25] in terms of visual fidelity, while maintaining comparable realism.
@@
  • We propose a novel dual-branch coding framework for extreme image compression. It decomposes features into general semantic and specific detail parts, dynamically adapting to the content variations of each individual image and enabling the flexible compression.
  • We propose a cross-branch interactive structure to eliminate the redundancy between the semantic and detail bitstreams, further facilitating compression efficiency.
  • DLF achieves SOTA performance on various datasets and metrics, delivering both high realism and fidelity reconstruction at extremely low bitrate.
@@ 2.1 Learned Image Compression
Variational autoencoder (VAE)-based image compression [34, 17, 27] has been widely studied. However, these methods, optimized for PSNR or MS-SSIM [42], often produce blurry image at low bitrates due to severe detail loss [5]. To enhance perceptual quality, Agustsson et al. [2] introduced the concept of generative image compression, using GANs [14] to generate the lost details. Mentzer et al. [32] proposed HiFiC, which features an advanced network structure and generative adversarial loss. The recent SOTA MS-ILLM [35] adopts a non-binary discriminator to greatly enhance statistical fidelity. Lee et al. [21] utilize text features to assist the encoder thus improving semantic consistency.
Despite these advancements, existing VAE-based generative image codecs [2, 32, 35, 21] are struggled to deliver pleasing quality at extremely low bitrates, such as 0.01 bpp. Consequently, recent work has begun exploring the use of generative visual tokenizers [12] and diffusion models [39] for extreme image compression.
@@ 2.2 Tokenizer-based Extreme Image Compression
Image tokenizers with discrete codebooks, such as VQVAE [41] and its subsequent work [38], can learn semantically rich tokens through latent clustering. The extracted tokens can be utilized for reconstruction, generation [12, 8], and other visual tasks [37, 45, 46]. The representative generative tokenizer VQGAN [12] introduces generative training and perceptual loss based on VQVAE, resulting in good perceptual quality. Several studies have employed such tokenizers to implement image codecs. Mao et al. [31] proposed adjusting the codebook size via K-means clustering in a pretrained VQGAN. By discarding and regenerating unimportant tokens via transformers, the bitrate is further reduced [43]. Jia et al. [16] introduced the transform coding in the latent space of VQGAN to achieve impressive performance. More recently, the 1-D tokenizer [47] improved semantic compression by reducing spatial redundancy; however, its fixed resolution poses challenges for flexible compression applications. In this paper, we reference the 1-D tokenizer for compact semantic compression in a patch-wise manner, enabling variable resolution. Furthermore, our work is tailored with adaptive detail coding, enhancing fidelity beyond that of the 1-D tokenizer alone.
Our work is related to HybridFlow [30], which combines discrete tokens from a VQGAN tokenizer with continuous features from an MLIC encoder [17]. However, our approach differs in three key aspects: (1) As adopting independent VQGAN tokenizer and MLIC encoder, HybridFlow contains non-trivial information redundancy between two branches. By contrast, DLF focuses on inter-branch redundancy reduction via cross-branch interaction. (2) HybridFlow relies on a manually pre-defined masking strategy for token reduction, rather than learning from the content, limiting its adaptability of diverse images. In comparison, DLF aggregates semantics into 1-D tokens through learning, thereby enhancing flexibility. (3) Since HybridFlow uses a fixed MLIC encoder, its detail features cannot adapt to generation errors in the tokenizer branch. By contrast, the detail encoder of DLF learns to dynamically correct generation through joint optimization, enhancing final quality. These innovations help DLF achieve 2.6×2.6$\times$2.6 × higher compression ratio than that of HybridFlow.
@@ 2.3 Diffusion-based Extreme Image Compression
Diffusion models [39, 36] have shown impressive generative capabilities. Text+Sketch [22] adopts Stable Diffusion [39] for image compression, conditioned on text prompts and structure maps. Despite improved realism, their fidelity remains unsatisfactory. In contrast, PerCo [7] fine-tunes a diffusion model within an image codec, incorporating vector-quantized spatial feature and global text description. It achieves remarkable reconstruction quality at extremely low bitrates below 0.01 bpp. Recently, DiffEIC [24] and RDEIC [25] use VAEs to compress the image latent for diffusion guidance to achieve superior performance. However, despite these guidance techniques, diffusion models still often generate irrelevant content at extremely low bitrates, leading to sub-optimal reconstruction fidelity. Compared with the diffusion-based methods, our proposed dual-branch coding framework delivers enhanced fidelity, competitive realism, and significantly faster decoding speed.
@@ 3.1 Dual-branch Coding Pipeline
Unlike existing methods [31, 43, 16] that employ the generative tokenizer [12] in a single branch, our DLF introduces an innovative dual-branch coding pipeline. As illustrated in Fig. 4, it decomposes the latent representation into semantic and detail components. In the semantic branch, DLF clusters high-level semantics into compact tokens by learning a codebook from a large training dataset. It then employs a detail branch to capture the diverse individual-level details through a large quantization space, complementing the semantic branch. In addition, DLF incorporates a cross-branch interaction mechanism to minimize redundancy between the semantic and detail branchs.
@@
For the semantic branch, we adopt a 1-D tokenizer [47] to learn dataset-level semantics clustering to facilitate efficient low-bitrate encoding. Specifically, we first partition $Emb(X)$ into $16\times 16$ windows. Within each window, both the $16\times 16$ 2-D image tokens and 32 extra 1-D tokens are fed into a vision transformer (ViT). Through cascaded attention blocks, these 1-D tokens efficiently aggregate essential semantic information from the image tokens, thereby minimizing spatial redundancy. Consequently, we encode these compact 1-D tokens to yield $y_{s}\in\mathbb{R}^{N\times C\times 32}$, where $N=\frac{h\times w}{16\times 16}$ denotes the number of windows in feature map.
@@
Quantization. We learn a codebook to encode the semantic latents $y_{s}$ through vector quantization (VQ). The resulting codebook indices are encoded using fixed-length coding for simplicity. For detail latents $y_{d}$, we employ scalar quantization (SQ) and use entropy model to estimate their distribution for arithmetic coding. Compared to VQ, SQ effectively leverages a much larger quantization space [33] to represent various individual-level details. In particular, we can easily apply different learnable quantization steps to SQ for spatially different regions. So the detail branch can adaptively allocate more bits to encode the most critical information. The quantization process in DLF can be formulated as:
@@
Fig. 5 presents the entropy distribution of the quantized detail features. For relatively common content, such as mountains and trees, the semantic branch can effectively encode them because the codebook is enriched with this type of information frequently occurring in the dataset. As a results, the detail branch just requires very low bits to represent these regions, even if they may contain high-frequency information. By contrast, relatively unique objects like the outlines of the castle and boat are more challenging to represent by highly clustered tokens, so the detail branch adaptively allocates more bits to encode these objects. This illustrates how DLF leverages the semantic and detail branches to dynamically encode images thus achieving both lower bitrate and improved visual fidelity.
Decoding. The quantized semantic and detail latents are decoded by their corresponding branch, resulting in the intermediate features $h_{s}\in\mathbb{R}^{C\times h\times w}$ and $h_{d}\in\mathbb{R}^{C\times h\times w}$:
@@
Subsequently, the intermediate features are fused through a latent adaptor and ultimately converted into reconstructed images $\hat{X}$ using a pixel generator $G$:
@@
Figure 6: (a) Interactive Transform (IT) module. For each window in semantic branch, we have $16\times 16$ feature tokens and 32 additional tokens, thus the length of the sequence $f_{s}$ is 288. (b) Detail transform block. We employs shifted window multi-head self-attention (W-MSA and SW-MSA) [28] and ConvNeXT block [29] in this module. (c) Semantic transform block. We adopts standard ViT module to extract semantics for each window.
@@ 3.2 Cross-branch Interactive Design
The dual-branch design of DLF requires eliminating redundancy between the semantic and detail bottlenecks. Furthermore, the window-wise encoding in the semantic branch requires inter-window interaction to capture global semantic information. To meet these requirements, we propose a cross-branch interaction design within DLF, implemented through a stacked Interactive Transform (IT) module. As depicted in Fig.6 (a), the detail feature $f_{d}\in\mathbb{R}^{c\times h\times w}$ is initially partitioned into $\tilde{f_{d}}\in\mathbb{R}^{N\times C\times 256}$ using the same strategy as in the semantic branch. Subsequently, the semantic and detail features are jointly fed into the multi-head self-attention layer to facilitate interaction. Finally, the processed detail feature is reshaped back into its original form.
Such design offers two significant advantages. First, the self-attention layer dynamically aggregates the semantic and detail information and reallocates them to each branch. Through this gradual aggregation and reallocation, the information redundancy between two branches can be effectively eliminated. Second, the interaction provides the semantic branch with cross-window perception. The detailed features contribute global information through this interaction, extending the semantic branch’s receptive field beyond individual windows and thereby enhancing semantic capturing capability.
@@
Figure 9: Qualitative examples on the CLIC2020 (full resolution).

### v3

@@ Abstract
Recent studies in extreme image compression have achieved remarkable performance by compressing the tokens from generative tokenizers. However, these methods often prioritize clustering common semantics within the dataset, while overlooking the diverse details of individual objects. Consequently, this results in suboptimal reconstruction fidelity, especially at low bitrates. To address this issue, we introduce a Dual-generative Latent Fusion (DLF) paradigm. DLF decomposes the latent into semantic and detail elements, compressing them through two distinct branches. The semantic branch clusters high-level information into compact tokens, while the detail branch encodes perceptually critical details to enhance overall fidelity. Additionally, we propose a cross-branch interactive design to reduce redundancy between the two branches, thereby minimizing the overall bit cost. Experimental results demonstrate the impressive reconstruction quality of DLF even below 0.01 bits per pixel (bpp). On the CLIC2020 test set, our method achieves bitrate savings of up to 27.93% on LPIPS and 53.55% on DISTS compared to MS-ILLM. Furthermore, DLF surpasses recent diffusion-based codecs in visual fidelity while maintaining a comparable level of generative realism. Project: https://dlfcodec.github.io/
@@ 1 Introduction
As image data continues to grow, achieving efficient compression while maintaining human perceptual quality has become crucial. At low bitrates, both traditional codecs [7, 11] and learning-based codecs [10, 18, 29] (optimized for PSNR) often produce blurry reconstructions due to substantial information degradation [6]. In this context, generative codecs [3, 34, 36], which enhance subjective quality by adding high-frequency details to reconstructed images, have attracted considerable attention. To achieve effective generative coding at extremely low bitrates, recent studies [33, 48, 17, 38] have employed generative visual tokenizers [43, 13] as foundational components. These tokenizers cluster common semantic content across the entire dataset into a compact codebook while simultaneously learning a powerful pixel generator based on the clustered information. As illustrated in Fig. 1 (a), if you do not look closely, many common semantic content within the dataset can be reconstructed with acceptable quality using a small codebook. Consequently, such codecs [33, 48, 17] only need to compress and transmit a few token indices, thereby significantly enhancing the compression ratio.
@@
However, when we zoom in each image to focus on the details of individual objects, we observe that the limited codebook capacity fails to capture these details accurately when compared with the ground truth. This inadequacy is particularly unfavorable for objects with precise geometric features, leading to a noticeable degradation in fidelity even without a reference for comparison. As illustrated in Fig. 1 (a), a reduced number of clusters results in severe distortion on a dog’s face. This observation underscores the limitation of the aforementioned codecs [33, 48, 17]: their backend tokenizer employs a single codebook to learn uniform latent clusters for all image content, failing to balance the demand between the common semantic learning across the dataset and the diverse detail representation of the individual objects. Therefore, the question arises: How can we compress the latent in a more flexible manner, dynamically adapting to the content variations of each image?
@@
To tackle this challenge, we introduce a novel extreme image coding paradigm named Dual-generative Latent Fusion (DLF). As illustrated in Fig. 3, DLF decomposes the image latent into semantic and detail components, employing two parallel branches to compress each component according to its demands. The semantic branch inherits the clustering capability of the generative tokenizer, capturing the common content at the dataset-level with few bits. Meanwhile, the detail branch complements the shortcomings of the single codebook tokenizer, representing the diverse details through a large quantization space. By training with a rate-distortion loss, the detail branch learns to adaptively allocate more bits to the most critical regions for enhanced fidelity, while avoiding the excessive data overhead. Finally, the decoded semantic and detail features are fused using a latent adaptor, aligning them into the same latent space for high-quality reconstruction using the pixel generator. As shown in Fig. 1 (b), our DLF not only reduces bit cost but also significantly enhances the detail fidelity while preserving the overall perceptual quality.
@@
By adopting a dual-branch coding design, DLF generates two bitstreams, similar to HybridFlow [32]. To minimize the overall bit cost, it is crucial to reduce redundancy between the two bitstreams. To this end, we introduce a cross-branch interaction mechanism that employs multiple attention-based layers to jointly process the intermediate features of both branches. This interaction not only optimizes bit allocation between the bottlenecks but also enables each branch to retain only the complementary information, thereby considerably reducing redundancy. By contrast, HybridFlow adopts two independent encoders without joint design, increasing bit cost due to redundancy.
@@
Powered by our advanced designs, DLF achieves state-of-the-art (SOTA) compression performance across various datasets and metrics, especially at extremely low bitrates such as 0.01 bpp. Qualitative results in Fig. 2 highlight the superiority of our method in terms of both generative realism and reconstruction fidelity. Compared to MS-ILLM [36], our approach achieves bitrate savings of 27.93% for LPIPS and 53.55% for DISTS on the CLIC2020 dataset. Additionally, our DLF significantly surpasses recent diffusion-based codecs [8, 26, 27] in terms of visual fidelity while maintaining comparable realism.
@@
  • We propose a novel dual-branch coding framework for extreme image compression. It decomposes features into general semantic and specific detail parts, dynamically adapting to the content variations of each individual image and enabling flexible compression.
  • We propose a cross-branch interactive structure to eliminate redundancy between the semantic and detail bitstreams, further facilitating compression efficiency.
  • DLF achieves SOTA performance on various datasets and metrics, delivering both high realism and faithful reconstruction at extremely low bitrates.
@@ 2.1 Learned Image Compression
Variational autoencoder (VAE)-based image codecs [18, 29, 44, 19, 56, 57] have been widely studied. However, these methods, mainly optimized for PSNR, produce blurry images at low bitrates due to severe detail loss [6]. To enhance perceptual quality, Agustsson et al. [3] introduced the concept of generative image compression, using GANs [15] to generate the lost details. Mentzer et al. [34] proposed HiFiC, which features an advanced network and a generative adversarial loss. The recent MS-ILLM [36] adopts a non-binary discriminator to greatly enhance statistical fidelity. Lee et al.[22] further utilize text features to assist the encoder, thus improving semantic consistency. Despite these advancements, existing VAE-based generative image codecs [3, 34, 36, 22] struggle to deliver pleasing quality at extremely low bitrates. Consequently, recent work has begun exploring the use of generative visual tokenizers [13] and diffusion models [41] for extreme image compression.
@@ 2.2 Tokenizer-based Extreme Image Compression
Image tokenizers with discrete codebooks, such as VQVAE [43] and its subsequent work [40], can learn semantically rich tokens through latent clustering. The extracted tokens can be utilized for reconstruction, generation [13, 9], and other visual tasks [39, 51, 52]. The representative generative tokenizer VQGAN [13] introduces generative training and perceptual loss based on VQVAE, resulting in enhanced perceptual quality. Several studies have employed such tokenizers to implement image codecs. Mao et al. [33] proposed adjusting the codebook size via K-means clustering in a pretrained VQGAN. By discarding and regenerating unimportant tokens via transformers, the bitrate is further reduced [48]. Jia et al. [17] introduced transform coding in the latent space of VQGAN to achieve impressive performance. More recently, the 1-D tokenizer [53] improved semantic compression by reducing spatial redundancy; however, its fixed resolution poses challenges for flexible compression applications. In this paper, we reference the 1-D tokenizer for compact semantic compression in a patch-wise manner, enabling variable resolution. Furthermore, our work is tailored with adaptive detail coding, enhancing fidelity beyond that of the 1-D tokenizer alone.
Our work is related to HybridFlow [32], which combines discrete tokens from a VQGAN tokenizer with continuous features from an MLIC encoder [18]. However, our approach differs in three key aspects: (1) By adopting independent VQGAN tokenizer and MLIC encoder, HybridFlow contains non-trivial information redundancy between the two branches. By contrast, DLF focuses on inter-branch redundancy reduction via cross-branch interaction. (2) HybridFlow relies on a manually pre-defined masking strategy for token reduction, rather than learning from the content, limiting its adaptability to the diverse images. In comparison, DLF aggregates semantics into 1-D tokens through learning, thereby enhancing flexibility. (3) Since HybridFlow uses a fixed MLIC encoder, its detail features cannot adapt to generation errors in the tokenizer branch. By contrast, the detail encoder of DLF learns to dynamically correct generation through joint optimization, enhancing final quality. These innovations help DLF achieve $2.6\times$ higher compression ratio than that of HybridFlow.
@@ 2.3 Diffusion-based Extreme Image Compression
Diffusion models [41, 37] have shown impressive capabilities in generative compression tasks [23, 8, 26, 27, 46, 55, 47, 49]. Text+Sketch [23] adopts Stable Diffusion [41] for image compression, conditioned on text prompts and structure maps. Despite improved realism, their fidelity remains unsatisfactory. In contrast, PerCo [8] fine-tunes a diffusion model within an image codec, incorporating vector-quantized spatial features and global text descriptions. It achieves remarkable reconstruction quality at extremely low bitrates below 0.01 bpp. Recently, DiffEIC [26] and RDEIC [27] use VAEs to compress the image latent for diffusion guidance to achieve superior performance. However, despite these guidance techniques, diffusion models still often generate irrelevant content at extremely low bitrates, leading to suboptimal reconstruction fidelity. Compared with diffusion-based methods, our proposed dual-branch coding framework delivers enhanced fidelity, competitive realism, and significantly faster decoding speed.
@@ 3.1 Dual-branch Coding Pipeline
Unlike existing methods [33, 48, 17] that employ the generative tokenizer [13] in a single branch, our DLF introduces an innovative dual-branch coding pipeline. As illustrated in Fig. 4, it decomposes the latent representation into semantic and detail components. In the semantic branch, DLF clusters high-level semantics into compact tokens by learning a codebook from a large training dataset. It then employs a detail branch to capture diverse individual-level details through a large quantization space, complementing the semantic branch. In addition, DLF incorporates a cross-branch interaction mechanism to minimize redundancy between the semantic and detail branches.
@@
For the semantic branch, we adopt a 1-D tokenizer [53] to learn dataset-level semantic clustering to facilitate efficient low-bitrate encoding. Specifically, we first partition $Emb(X)$ into $16\times 16$ windows. Within each window, both the $16\times 16$ 2-D image tokens and 32 extra 1-D tokens are fed into a vision transformer (ViT). Through cascaded attention blocks, these 1-D tokens efficiently aggregate essential semantic information from the image tokens, thereby minimizing spatial redundancy. Unlike previous methods [48, 32] that use manually pre-defined masks for token reduction, our approach compresses semantics into fewer tokens through large-scale learning, offering greater flexibility in handling various images. Consequently, we encode these compact 1-D tokens to yield $y_{s}\in\mathbb{R}^{N\times C\times 32}$, where $N=\frac{h\times w}{16\times 16}$ denotes the number of windows in feature map.
@@
Quantization. We learn a codebook to encode the semantic latents $y_{s}$ through vector quantization (VQ). The resulting codebook indices are encoded using fixed-length coding for simplicity. For detail latents $y_{d}$, we employ scalar quantization (SQ) and use quadtree partition-based entropy model [24, 25] to estimate their distribution for arithmetic coding. Compared to VQ, SQ effectively leverages a much larger quantization space [35] to represent various individual-level details. In particular, we can easily apply different learnable quantization steps to SQ for spatially varying regions. So the detail branch can adaptively allocate more bits to encode the most critical information. The quantization process in DLF can be formulated as:
@@
Figure 6: (a) Interactive Transform (IT) module. For each window in semantic branch, we have $16\times 16$ feature tokens and 32 additional tokens, thus the length of the sequence $f_{s}$ is 288. (b) Detail transform block. We employs shifted window based multi-head self-attention (W-MSA and SW-MSA) [30] and ConvNeXT block [31] in this module. (c) Semantic transform block. We adopts standard ViT module to extract semantics for each window.
Fig. 5 presents the entropy distribution of the quantized detail features. For relatively common content, such as mountains and trees, the semantic branch can effectively encode them because the codebook is enriched with this type of information frequently occurring in the dataset. Consequently, the detail branch requires very few bits to represent these regions, even if they may contain high-frequency information. By contrast, relatively unique objects like the outlines of the castle and boat are more challenging to represent with highly clustered tokens, so the detail branch adaptively allocates more bits to encode these objects. This illustrates how DLF leverages the semantic and detail branches to dynamically encode images, thus achieving both lower bitrate and improved visual fidelity.
Decoding. The quantized semantic and detail latents are decoded by their corresponding branch, producing the intermediate features $h_{s}\in\mathbb{R}^{C\times h\times w}$ and $h_{d}\in\mathbb{R}^{C\times h\times w}$:
@@
Subsequently, the intermediate features are fused through a latent adaptor and ultimately converted into the reconstructed image $\hat{X}$ using a pixel generator $G$:
@@ 3.2 Cross-branch Interactive Design
The dual-branch design of DLF requires eliminating redundancy between the semantic and detail bottlenecks. Furthermore, the window-wise encoding in the semantic branch requires inter-window interactions to capture global semantic information. To meet these requirements, we propose a cross-branch interaction design within DLF, implemented through a stacked Interactive Transform (IT) module. As depicted in Fig. 6 (a), the detail feature $f_{d}\in\mathbb{R}^{c\times h\times w}$ is initially partitioned into $\tilde{f_{d}}\in\mathbb{R}^{N\times C\times 256}$ using the same strategy as in the semantic branch. Subsequently, the semantic and detail features are jointly fed into the multi-head self-attention layer to facilitate interaction. Finally, the processed detail feature is reshaped back into its original form.
Such design offers two significant advantages. First, the self-attention layer dynamically aggregates the semantic and detail information through joint training, reallocating them to each branch. This gradual aggregation and reallocation not only mitigate redundancy between branches but also enable detail information to dynamically correct generation errors in the semantic branch. Second, the interaction provides the semantic branch with cross-window perception. The detail feature contributes global information through this interaction, extending the semantic branch’s receptive field beyond individual windows and thereby enhancing semantic capturing capability.
@@
Figure 9: Qualitative examples on the CLIC2020 ($768\times 768$) dataset. More comparisons are in supplementary.